what is the velocity of the run when a student runs 2 laps around a track in 120 seconds when the distance of the track is 400
1
Expert's answer
2012-10-31T08:37:08-0400
The student make 2 laps around track of length 400 m, soit runs 2*400=800 m.
It takes the time 120s=2minutes, whence its velocity is
v = 800 m/ 2minutes = 400 m/minute = 400 m / 60 s = 6.67 m/s